OpenStudy (anonymous):

pleasee Helpppp copy and complete this sequence of calculations(^2 means squared - a) 3^2 - 1^2 = 4^2 - 2^2 = 5^2 - 3^2 = 6^2 - 4^2 = B) describe the sequence of answers to the calculations C) describe a rule to find the answer without squaring the numbers. D) use the rule to find the answer to 15 squared - 13 squared

OpenStudy (anonymous):

square each first, then subtract second

OpenStudy (anonymous):

for example the last one is \(6^2 - 4^2 \) so first compute \(6\times 6=36\) and \(4\times 4=16\) and then compute \[36-16=20\]

OpenStudy (anonymous):

i think that you are also supposed to notice that each \[a^2-b^2\] is the same as \[(a+b)(a-b)\] for example \[6^2-4^2=36-16=20=(6+4)(6-4)=10\times 2\]

OpenStudy (anonymous):

I've done the first bit but i need b c and d /: xxx